I need a 9% solution of Sulfuric acid. The base acid is 37% by volume and I need 55 gallons.
Answer by Theo(13342)

you need a 9% solution.
your base solution is 37% acid.
x = your base solution.
y = a 0% solution that I'm assuming you're need to mix it with if that's possible.
your equations are:
x + y = 55
.37x + 0y = .09*55
simplify these equations to get:
x + y = 55
.37x = 4.95
divide both sides of .37x = 4.95 to get x = 4.95 / .37 which becomes x = 13.37837838.
since x + y = 55, this means that y has to be equal to 55 - 13.37837838 = 41.62162162

you have to mix 41.62162162 gallons of 0% acid with 13.37837838 gallons of 37% acid to get 55 gallons of 9% acid.

let's go the math and see if that works.
.37 * 13.37837838 + 0 * 41.62162162 = .09 * 55
simplify to get:
4.95 + 0 = 4.95

answer is confirmed to be correct.

you need to mix 13.37837838 gallons of 37% acid solution with 41.62162162 gallons of 0% acid solution to get 55 gallons of 9% acid solution.